T.C. = 3,000 + 24 Q - 4 Q² + 0.4 Q³  

1. If the market price is $200 per unit, what would this firm’s optimum level of production and sales (Q) be per week? Please round-off your Q to the nearest unit.

2. What is the total profit at optimum Q? Would this firm continue operations? Why or why not?

3. Below what market price would it be preferable for this firm to shut-down? Why?
Answer #1

1 TC= 3000l 240- 40 2+0 y Q3 Mc dlag(t) = d cao (3000 +2uco a Qu_80 + 1.20² to (3000 + suo 4% to uQ3 P = 200 At equilibrium iAt this price, VC = TC - TFC = $3976.23-$3000 = $976.23

SINCE THE VARIABLE COST IS LESSER THAN THE TOTAL REVENUE THE FORM WILL STILL CONTINUE TO THE OPERATIONS

3) AVC = VC/Q = 976.23/15.89 = 61.44

Since the the the average variable cost is less than price by a huge margin till the point the price decreases to the average variable cost the firm will continue